OpenStudy (anonymous):

Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. a. passes through (-1, 4) and (-5, 2)

OpenStudy (karatechopper):

point slope form cuz its easy for my to work with y-4=1/3(x+1) y-4=1/3x+1/3 (y-4=1/3x+1/3)3 3y-4=x+1 3y-4+4=x+1-4 3y=x-3 3y/3=x-3/3 y=1/3x-1